# Local Differential Privacy for Distributed Stochastic Aggregative Optimization with Guaranteed Optimality

### v2

@@ VII-A Auxiliary lemmas
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]$ (69)
$\textstyle=\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\tilde{W}\otimes I_{r})^{t}\Xi_{y}^{1}\rangle]$
$\textstyle\quad+\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$
$\textstyle\quad+\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Xi_{g}^{p}\rangle].$
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\tilde{W}\otimes I_{r})^{t}\Xi_{y}^{1}\rangle]$ (70)
$\textstyle\quad\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{4d_{0}}{\kappa(t+1)^{2-2v_{y}}},$
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$ (71)
$\textstyle=\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\!\boldsymbol{\bar{x}}^{t-1},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$
$\textstyle+\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t-1}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\!\otimes\!I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle].$
Since $\boldsymbol{\bar{x}}^{t-1}-\boldsymbol{x}^{*}$ and $\Xi_{\chi}^{t}$ are independent, we have $\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t-1}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\tilde{W}\!\otimes\!I_{r})^{0}K_{1}\Xi_{\chi}^{t}\rangle]=0$. Hence, the second term on the right hand side of (71) can be rewritten as
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t-1}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\!\otimes\!I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$ (72)
$\textstyle=\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t-1}\!-\!\boldsymbol{x}^{*},$
@@
Applying an argument similar to the derivation of (72), we iterate (71) from 00 to $t$ to obtain
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$ (73)
$\textstyle=\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{\bar{x}}^{t-1},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$
$\textstyle\quad+\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t-1}-\boldsymbol{\bar{x}}^{t-2},$
@@
[…]
$\textstyle\quad+\mathbb{E}[\langle\boldsymbol{\bar{x}}^{1}-\boldsymbol{\bar{x}}^{0},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\tilde{W}\otimes I_{r})^{t-1}K_{1}\Xi_{\chi}^{1}\rangle],$
where we used $\mathbb{E}[\langle\boldsymbol{\bar{x}}^{0}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\tilde{W}\!\otimes\!I_{r})^{t-1}K_{1}\Xi_{\chi}^{1}\rangle]=0$ based on $\boldsymbol{\bar{x}}^{0}-\boldsymbol{x}^{*}$ and $\Xi_{\chi}^{1}$ being independent.
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{\bar{x}}^{t-1},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$ (74)
$\textstyle\leq\sqrt{\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{\bar{x}}^{t-1}\|^{2}]}$
×$\textstyle\quad\times\frac{1}{\lambda_{y}^{t}}\sqrt{\mathbb{E}[\|M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\|^{2}]}.$
By defining $\boldsymbol{\bar{x}}^{t}=\mathrm{col}(\bar{x}_{(1)}^{t},\cdots,\bar{x}_{(m)}^{t})\in\mathbb{R}^{n}$ and using the dynamics of $x_{i}^{t}$ in Algorithm 1, we have
$\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{\bar{x}}^{t-1}\|^{2}]\leq n(\sigma_{x}^{t-1})^{2}+\frac{(\lambda_{x}^{t-1})^{2}}{m^{2}}\mathbb{E}[\|\boldsymbol{u}^{t-1}\|^{2}].$ (75)
@@
$\textstyle\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{\bar{x}}^{t-1}\|^{2}]\leq\frac{n\sigma_{x}^{2}}{t^{2\varsigma_{x}}}+\frac{C_{z}c_{u1}(\lambda_{x}^{0})^{2}}{m^{2}(\lambda_{z}^{0})^{2}t^{\beta_{z}+2v_{x}-2v_{z}}}$ (76)
@@
with $d_{1}=n\sigma_{x}^{2}+m^{-1}(\lambda_{x}^{0})^{2}(\lambda_{z}^{0})^{-2}(c_{u1}(C_{z}+r\sigma_{z}^{2})+c_{u2})$. Here, in the derivation, we have used the definition $\beta_{z}=\min\{2v_{z}+2\varsigma_{y}-2v_{y},2\varsigma_{z}\}$ and the relation $v_{x}-v_{y}>v_{x}-v_{z}>\varsigma_{x}$.
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{\bar{x}}^{t-1},-\frac{M_{\boldsymbol{\tau}^{t}}}{\lambda_{y}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]\leq\frac{\sqrt{2^{2\varsigma_{x}}d_{1}d_{2}}}{\lambda_{y}^{0}(t+1)^{\beta_{2}}},$
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$ (79)
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Xi_{g}^{p}\rangle]$ (80)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{\bar{L}_{f}^{2}\sum_{p=1}^{t}((1-|\rho_{2}|)^{2})^{t-p}\mathbb{E}[\|\Xi_{g}^{p}\|^{2}]}{\kappa(\lambda_{y}^{t})^{2}}.$
@@
$\textstyle\mathbb{E}[\|\Xi_{g}^{p}\|^{2}]\leq\frac{4C_{x}L_{g}^{2}(\lambda_{y}^{0})^{2}}{p^{2v_{y}+2\varsigma_{x}}}+\frac{4C_{z}L_{g}^{2}c_{u1}(\lambda_{y}^{0})^{2}(\lambda_{x}^{0})^{2}}{(\lambda_{z}^{0})^{2}p^{\beta_{z}+2v_{y}+2v_{x}-2v_{z}}}$
@@
where $\alpha_{1}$ is given by $\alpha_{1}=\min\{2v_{x}+2\varsigma_{y},2v_{y}+1,2v_{y}+2\varsigma_{x}\}$ and $d_{4}=4L_{g}^{2}(\lambda_{x}^{0})^{2}(\lambda_{y}^{0})^{2}(\lambda_{z}^{0})^{-2}(C_{z}c_{u1}+mrc_{u1}\sigma_{z}^{2})+2L_{g}^{2}(\lambda_{y}^{0})^{2}(2C_{x}+n\sigma_{x}^{2}+2mc_{u2}(\lambda_{x}^{0})^{2})+2m(2+v_{y}^{2})(\lambda_{y}^{0})^{2}\sigma_{g,0}^{2}+2md_{g}^{2}v_{y}^{2}(\lambda_{y}^{0})^{2}$.
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Xi_{g}^{p}\rangle]$ (82)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{5}}{\kappa(t+1)^{\beta_{1}}},$
@@
By substituting (70), (79), (82) into (69) and using the relation $12-2v_{y}>$ 1, we arrive at (61). ∎
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},\nabla\boldsymbol{g}(\boldsymbol{x}^{t})(\boldsymbol{1}\!\otimes\!\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}\!\otimes\!g(\boldsymbol{x}^{t})))\!-\!\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\boldsymbol{\tilde{z}}^{t}\rangle]$
$\textstyle\leq\frac{(7+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]\!+\!\frac{b_{3}}{\kappa(t+1)^{\beta_{1}}}\!+\!\frac{b_{4}}{(t+1)^{\beta_{3}}},$ (83)
@@
$\textstyle\!d_{6}=2C_{x}+\frac{2c_{u1}(\lambda_{x}^{0})^{2}(C_{z}+mr\sigma_{z}^{2})}{(\lambda_{z}^{0})^{2}}+2mc_{u2}(\lambda_{x}^{0})^{2}+n\sigma_{x}^{2};$
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},(\nabla\boldsymbol{g}(\boldsymbol{x}^{t})\!-\!\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t}))$ (86)
×$\textstyle\quad\times(\boldsymbol{1}_{m}\!\otimes\!\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t})))\rangle]$
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{m\sigma_{g,1}^{2}L_{f}^{2}}{\kappa(t+1)}.$
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})(\boldsymbol{1}_{m}\otimes\nabla_{y}\bar{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t})))$ (87)
@@
$\textstyle=\mathbb{E}[\langle\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})^{\top}(\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}),$
@@
To characterize the right hand side of (87), we first estimate an upper bound on $\mathbb{E}[\|\Xi_{x}^{t}-\Xi_{x}^{t-1}\|^{2}]$ with denoting $\Xi_{x}^{t}\triangleq\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})^{\top}(\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*})$. We have the following decomposition:
$\textstyle\mathbb{E}[\|\Xi_{x}^{t}-\Xi_{x}^{t-1}\|^{2}]=\mathbb{E}[\|\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})^{\top}(\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*})$
$\textstyle\quad-\nabla\boldsymbol{g}^{t-1}(\boldsymbol{x}^{t-1})^{\top}(\boldsymbol{\bar{x}}^{t-1}-\boldsymbol{x}^{*})\|^{2}]$
$\textstyle\leq\mathbb{E}[\|(\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})-\nabla\boldsymbol{g}(\boldsymbol{x}^{t}))^{\top}(\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*})\|^{2}]$
$\textstyle\quad+\mathbb{E}[\|(\nabla\boldsymbol{g}^{t-1}(\boldsymbol{x}^{t-1})-\nabla\boldsymbol{g}(\boldsymbol{x}^{t-1}))^{\top}(\boldsymbol{\bar{x}}^{t-1}-\boldsymbol{x}^{*})\|^{2}]$
$\textstyle\quad+\mathbb{E}[\|\nabla\boldsymbol{g}(\boldsymbol{x}^{t})^{\top}(\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*})\!-\!\nabla\boldsymbol{g}(\boldsymbol{x}^{t-1})^{\top}(\boldsymbol{\bar{x}}^{t-1}\!-\!\boldsymbol{x}^{*})\|^{2}]$

### v3

@@ VII-A Auxiliary lemmas
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]$ (63)
$\textstyle=\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\tilde{W}\otimes I_{r})^{t}\Xi_{y}^{1}\rangle]$
$\textstyle\quad+\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$
$\textstyle\quad+\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Xi_{g}^{p}\rangle].$
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\tilde{W}\otimes I_{r})^{t}\Xi_{y}^{1}\rangle]$ (64)
$\textstyle\quad\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{4d_{0}}{\kappa(t+1)^{2-2v_{y}}},$
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$ (65)
$\textstyle=\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\!\mathbf{\bar{x}}^{t-1},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$
$\textstyle+\mathbb{E}[\langle\mathbf{\bar{x}}^{t-1}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\!\otimes\!I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle].$
Since $\mathbf{\bar{x}}^{t-1}-\boldsymbol{x}^{*}$ and $\Xi_{\chi}^{t}$ are independent, we have $\mathbb{E}[\langle\mathbf{\bar{x}}^{t-1}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\tilde{W}\!\otimes\!I_{r})^{0}K_{1}\Xi_{\chi}^{t}\rangle]=0$. Hence, the second term on the right hand side of (65) can be rewritten as
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t-1}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\!\otimes\!I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$ (66)
$\textstyle=\mathbb{E}[\langle\mathbf{\bar{x}}^{t-1}\!-\!\boldsymbol{x}^{*},$
@@
Applying an argument similar to the derivation of (66), we iterate (65) from 0 to $t$ to obtain
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$ (67)
$\textstyle=\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\mathbf{\bar{x}}^{t-1},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$
$\textstyle\quad+\mathbb{E}[\langle\mathbf{\bar{x}}^{t-1}-\mathbf{\bar{x}}^{t-2},$
@@
$\textstyle\quad$$\vdots$
$\textstyle\quad+\mathbb{E}[\langle\mathbf{\bar{x}}^{1}-\mathbf{\bar{x}}^{0},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\tilde{W}\otimes I_{r})^{t-1}K_{1}\Xi_{\chi}^{1}\rangle],$
where we used $\mathbb{E}[\langle\mathbf{\bar{x}}^{0}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\tilde{W}\!\otimes\!I_{r})^{t-1}K_{1}\Xi_{\chi}^{1}\rangle]=0$ based on $\mathbf{\bar{x}}^{0}-\boldsymbol{x}^{*}$ and $\Xi_{\chi}^{1}$ being independent.
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\mathbf{\bar{x}}^{t-1},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$ (68)
$\textstyle\leq\sqrt{\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\mathbf{\bar{x}}^{t-1}\|^{2}]}$
$\textstyle\quad\times\frac{1}{\lambda_{y}^{t}}\sqrt{\mathbb{E}[\|M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\|^{2}]}.$
By defining $\mathbf{\bar{x}}^{t}=\mathrm{col}(\bar{x}_{(1)}^{t},\cdots,\bar{x}_{(m)}^{t})\in\mathbb{R}^{n}$ and using the dynamics of $x_{i}^{t}$ in Algorithm 1, we have
$\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\mathbf{\bar{x}}^{t-1}\|^{2}]\leq n(\sigma_{x}^{t-1})^{2}+\frac{(\lambda_{x}^{t-1})^{2}}{m^{2}}\mathbb{E}[\|\boldsymbol{u}^{t-1}\|^{2}].$ (69)
@@
$\textstyle\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\mathbf{\bar{x}}^{t-1}\|^{2}]\leq\frac{n\sigma_{x}^{2}}{t^{2\varsigma_{x}}}+\frac{\bar{C}_{z}c_{u1}(\lambda_{x}^{0})^{2}}{m^{2}(\lambda_{z}^{0})^{2}t^{\beta_{z}+2v_{x}-2v_{z}}}$ (70)
@@
with $d_{1}=n\sigma_{x}^{2}+m^{-1}(\lambda_{x}^{0})^{2}(\lambda_{z}^{0})^{-2}(c_{u1}(\bar{C}_{z}+r\sigma_{z}^{2})+c_{u2})$. Here, in the derivation, we have used the definition $\beta_{z}=\min\{2v_{z}+2\varsigma_{y}-2v_{y},2\varsigma_{z}\}$ and the relation $v_{x}-v_{y}>v_{x}-v_{z}>\varsigma_{x}$.
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}\!-\!\mathbf{\bar{x}}^{t-1},-\frac{M_{\boldsymbol{\tau}^{t}}}{\lambda_{y}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]\leq\frac{\sqrt{2^{2\varsigma_{x}}d_{1}d_{2}}}{\lambda_{y}^{0}(t+1)^{\beta_{2}}},$
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\chi}^{p}\rangle]$ (73)
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Xi_{g}^{p}\rangle]$ (74)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{\bar{L}_{f}^{2}\sum_{p=1}^{t}((1-|\rho_{2}|)^{2})^{t-p}\mathbb{E}[\|\Xi_{g}^{p}\|^{2}]}{\kappa(\lambda_{y}^{t})^{2}}.$
@@
$\textstyle\mathbb{E}[\|\Xi_{g}^{p}\|^{2}]\leq\frac{4\bar{C}_{x}L_{g}^{2}(\lambda_{y}^{0})^{2}}{p^{2v_{y}+2\varsigma_{x}}}+\frac{4\bar{C}_{z}L_{g}^{2}c_{u1}(\lambda_{y}^{0})^{2}(\lambda_{x}^{0})^{2}}{(\lambda_{z}^{0})^{2}p^{\beta_{z}+2v_{y}+2v_{x}-2v_{z}}}$
@@
where $\alpha_{1}$ is given by $\alpha_{1}=\min\{2v_{x}+2\varsigma_{y},2v_{y}+1,2v_{y}+2\varsigma_{x}\}$ and $d_{4}=4L_{g}^{2}(\lambda_{x}^{0})^{2}(\lambda_{y}^{0})^{2}(\lambda_{z}^{0})^{-2}(\bar{C}_{z}c_{u1}+mrc_{u1}\sigma_{z}^{2})+2L_{g}^{2}(\lambda_{y}^{0})^{2}(2\bar{C}_{x}+n\sigma_{x}^{2}+2mc_{u2}(\lambda_{x}^{0})^{2})+2m(2+v_{y}^{2})(\lambda_{y}^{0})^{2}\sigma_{g,0}^{2}+2md_{g}^{2}v_{y}^{2}(\lambda_{y}^{0})^{2}$.
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Xi_{g}^{p}\rangle]$ (76)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{5}}{\kappa(t+1)^{\beta_{1}}},$
@@
By substituting (64), (73), (76) into (63) and using the relation $2-2v_{y}>1$, we arrive at (55). ∎
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},\nabla\boldsymbol{g}(\boldsymbol{x}^{t})(\boldsymbol{1}\!\otimes\!\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}\!\otimes\!g(\boldsymbol{x}^{t})))\!-\!\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\boldsymbol{\tilde{z}}^{t}\rangle]$
$\textstyle\leq\frac{(7+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]\!+\!\frac{b_{3}}{\kappa(t+1)^{\beta_{1}}}\!+\!\frac{b_{4}}{(t+1)^{\beta_{3}}},$ (77)
@@
$\textstyle\!d_{6}=2\bar{C}_{x}+\frac{2c_{u1}(\lambda_{x}^{0})^{2}(\bar{C}_{z}+mr\sigma_{z}^{2})}{(\lambda_{z}^{0})^{2}}+2mc_{u2}(\lambda_{x}^{0})^{2}+n\sigma_{x}^{2};$
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},(\nabla\boldsymbol{g}(\boldsymbol{x}^{t})\!-\!\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t}))$ (80)
$\textstyle\quad\times(\boldsymbol{1}_{m}\!\otimes\!\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t})))\rangle]$
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{m\sigma_{g,1}^{2}L_{f}^{2}}{\kappa(t+1)}.$
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})(\boldsymbol{1}_{m}\otimes\nabla_{y}\bar{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t})))$ (81)
@@
$\textstyle=\mathbb{E}[\langle\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})^{\top}(\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}),$
@@
To characterize the right hand side of (81), we first estimate an upper bound on $\mathbb{E}[\|\Xi_{x}^{t}-\Xi_{x}^{t-1}\|^{2}]$ with denoting $\Xi_{x}^{t}\triangleq\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})^{\top}(\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*})$. We have the following decomposition:
$\textstyle\mathbb{E}[\|\Xi_{x}^{t}-\Xi_{x}^{t-1}\|^{2}]=\mathbb{E}[\|\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})^{\top}(\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*})$
$\textstyle\quad-\nabla\boldsymbol{g}^{t-1}(\boldsymbol{x}^{t-1})^{\top}(\mathbf{\bar{x}}^{t-1}-\boldsymbol{x}^{*})\|^{2}]$